A capacitor when charged to 1000v, takes 2micro Coulumb of charge. Find the capacity of the capacitor.

Grade:11

1 Answers

Khushi goyal
5 years ago
Capacity= capacitance =C=Q/V = 2×10^-6÷1000 = 2 ×10^-9 F 
Hope it helps
Please approve  to get more help